Calculate the pH of each two-component solution: a. 0.265 M HNO3 and 0.102 M HCN b. 0.115M NaClO2 and 0.0500M NH3 c. 0.265M HNO2 and 0.350 M NaOH
Answer #1

a) 0.265 M HNO3 and 0.102 M HCN

HNO3 is strong acid and HCN is the weak acid. so pH depends on HNO3

pH = -log [H+]

pH = -log (0.265)

pH = 0.578

b)   0.115M NaClO2 and 0.0500M NH3

[OH-] = sqrt (Kb1 x C   + Kb2 x C )

           = sqrt (5.6 x 10^-4 x 0.115 + 1.8 x 10^-5 x 0.0500 )

            = 8.03 x 10^-3 M

pOH = 2.10

pH + pOH = 14

pH = 11.90

c)   0.265M HNO2 and 0.350 M NaOH

NaOH left = 0.350 -0.265

                = 0.085 M

pOH = 1.07

pH + pOH = 14

pH = 12.93